$SPAD/src/input ZLINDEP.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input generated from IntegerLinearDependenceXmpPage
)clear all
M := SQMATRIX(2,INT)
m1: M := squareMatrix matrix [[1, 2], [0, -1]]
m2: M := squareMatrix matrix [[2, 3], [1, -2]]
m3: M := squareMatrix matrix [[3, 4], [2, -3]]
linearlyDependentOverZ? vector [m1, m2, m3]
c := linearDependenceOverZ vector [m1, m2, m3]
c.1 * m1 + c.2 * m2 + c.3 * m3
solveLinearlyOverQ(vector [m1, m3], m2)

———-

References^{[1] nothing}

4